# Dynamic Benchmark Construction for Evaluating Large Language Models on Real-World Codes

### v1

@@ 4.1 Experimental Setup
Models Evaluated: We evaluated a diverse set of state-of-the-art Large Language Models, including both open-source and closed-source options. The complete list is provided in Table 3.
Evaluation Metric: We adopt Pass@1 [40] as our primary evaluation metric, widely used in code generation studies [19], which effectively reflects a developer’s experience with coding assistants.
Inference Settings: All models were evaluated in a zero-shot setting without example prompt-completion pairs. For Pass@1 evaluation, we used greedy decoding (temperature 0), following standard practices in deterministic code generation[23, 21, 8]. Open-source models with fewer than 32B parameters were served using vLLM [41], while larger open-source and closed-source models were accessed via official APIs.
Evaluation Process: Each model was evaluated in an automated, isolated environment. For each problem, the model received a generated instruction (including a refined docstring and function signature). The generated code was executed using language-specific Testcase Runners against PBT-generated test cases. Differential testing compared the outputs of the generated code with the ground-truth implementation for all test cases.
Table 2: Pass@1 Performance of LLMs on SC vs. WSC problems (Python implementation).
Model Name SC (%) WSC (%)
Claude-3.7-sonnet 39.9 49.5
Gemini-2.5-flash-preview 39.9 53.0
Doubao-1-5-pro-32k-250115 36.3 53.0
DeepSeek-V3 39.3 50.5
Llama-4-maverick 39.9 53.5
Qwen3-235b-a22b 37.8 50.0
Llama-4-scout 31.1 42.9
Llama-3.3-70b-Instruct 37.3 45.6
Qwen3-32b 35.2 45.5
Qwen2.5-Coder-32b-Instruct 36.2 44.4
Gemma-3-27b-it 34.2 46.0
Mistral-small-3.1-24b-Instruct 31.6 44.4
Qwen2.5-Coder-14b-Instruct 31.6 43.4
Qwen3-8b 28.5 39.9
Llama-3.2-3b-Instruct 18.1 23.2
Qwen3-1.7b 14.5 20.2
4.2 Performance Across Multiple Languages
To investigate the cross-domain implementation consistency of LLMs, we evaluate their performance on the SC subset of Code2Bench-2505 across five programming languages using the same instructions and test cases. This setup reveals how effectively LLMs translate universal task descriptions into diverse linguistic contexts.
Table 3: Pass@1 Performance of LLMs on Self-Contained (SC) problems across different languages.
Model Name Open Source Parameters Python (%) JS (%) TS (%) Java (%) Go (%)
Claude-3.7-sonnet No - 39.9 33.7 35.2 29.5 21.2
Gemini-2.5-flash-preview No - 39.9 35.2 34.1 28.5 27.5
Doubao-1-5-pro-32k-250115 No - 36.3 34.7 35.2 25.9 28.0
DeepSeek-V3 Yes 671B 39.3 34.2 32.6 25.9 23.8
Llama-4-maverick Yes 400B 39.4 34.2 30.6 22.8 11.4
Qwen3-235b-a22b Yes 235B 37.8 29.5 33.2 27.5 22.3
Llama-4-scout Yes 109B 31.1 25.4 28.0 25.4 6.2
Llama-3.3-70b-Instruct Yes 70B 37.3 27.5 27.5 24.4 15.5
Qwen3-32b Yes 32B 35.2 28.0 30.1 24.9 10.4
Qwen2.5-Coder-32b-Instruct Yes 32B 36.2 34.2 32.1 24.4 17.1
Gemma-3-27b-it Yes 27B 34.2 29.5 28.5 22.8 21.2
Mistral-small-3.1-24b-Instruct Yes 24B 31.6 30.6 23.8 22.7 13.0
Qwen2.5-Coder-14b-Instruct Yes 14B 31.6 29.5 29.5 22.8 16.6
Qwen3-8b Yes 8B 28.5 20.7 21.2 17.1 12.0
Llama-3.2-3b-Instruct Yes 3B 18.1 9.3 7.8 6.7 3.1
Qwen3-1.7b Yes 1.7B 14.5 10.4 9.3 9.3 1.6
Table 3 summarizes the Pass@1 results for each model. Performance varies significantly across languages, with Python consistently yielding the highest scores (e.g., Claude-3.7-sonnet achieving around 40% Pass@1). In contrast, models exhibit substantial drops in other languages (e.g., DeepSeek-V3 from 39.3% in Python to 23.8% in Go). These results highlight challenges in generalizing language-agnostic task descriptions, as well as language-specific hurdles such as Go’s complexity and TypeScript/Java’s strictness. Collectively, these findings underscore significant limitations in LLMs’ ability to achieve robust cross-lingual code generation on real-world codes.
4.3 Performance With Limited Dependencies
This subsection evaluates the impact of external dependencies on LLMs’ code generation performance for WSC tasks. By comparing SC and WSC tasks, we assess models’ ability to generalize to tasks involving library-specific data structures, and intricate logic on basic types. The analysis focuses on 16 models evaluated on 193 SC and 198 WSC problems in Python.
Table 2 presents the Pass@1 results for each model on SC and WSC subsets. Contrary to the common assumption that external dependencies increase task difficulty [23], our results reveal a surprising trend: models consistently achieve higher Pass@1 scores on WSC tasks compared to SC tasks in real-world scenarios. To explain this, we analyze the structural characteristics of the tasks. SC tasks have higher average cyclomatic complexity (5.21 vs. 2.23) and Lines of Code (LoC) (21.25 vs. 17.25) than WSC tasks. This suggests that the pure logical structure and control flow complexity within real-world SC functions tend to be more intricate than the core logic within real-world WSC functions. We hypothesize that LLMs, being strong pattern matchers, are relatively proficient at replicating common library interaction patterns found in WSC tasks, even if these involve complex data structures. Conversely, they struggle more with accurately reproducing the diverse, non-standard complex logic, intricate branching, and subtle edge cases often present in SC tasks, which require deeper algorithmic understanding rather than pattern recall. The detailed analysis of failure modes in Section 4.4 provides further qualitative evidence supporting this hypothesis.
4.4 Discussion and Analysis of Failure Modes
To gain deeper insights into why LLMs succeed or fail on different task types, we conducted a detailed analysis of the failure modes observed across all evaluated models. We classify outcomes into three categories: Execution Errors (code fails to run), Test Failures (code runs but produces incorrect output), and Success (code runs and passes all tests). Figure 3 presents the average distribution of these outcomes across all evaluated models for SC tasks in different programming languages and WSC tasks in Python. A key observation is the stark contrast in failure composition between SC tasks and WSC tasks. For SC tasks across all languages, the majority of failures are Test Failures, indicating that models are relatively successful at generating syntactically correct and runnable code, but frequently struggle with implementing the precise logic and handling edge cases. This pattern aligns with our hypothesis that the primary challenge in real-world SC tasks lies in their complex internal logic. In contrast, for WSC tasks in Python, the proportion of Execution Errors is significantly higher compared to SC tasks. This suggests that models are more prone to generating code that fails to run when interacting with external libraries, indicating issues with correct API usage, imports, or handling library-specific data structures. While Test Failures still occur in WSC, the increased rate of Execution Errors supports the hypothesis that handling the details of library interactions is a significant challenge. Furthermore, within SC tasks, we observe some variation across languages; for instance, Go shows a particularly high rate of Execution Errors, likely reflecting challenges related to its strictness. This analysis of failure modes provides strong evidence supporting our earlier findings regarding the differential challenges of SC and WSC tasks and offers diagnostic insights into the nature of LLM failures on real-world code. More details can be found in Appendix I.2.
4.5 Benchmarking Real-World Complexities: Code2Bench-2505 vs. EvalPlus
Figure 2: Average Performance and Failure Type Distribution across Task Types (SC vs WSC).
Figure 3: Comparing Pass@1 on Code2Bench-2505 (Python SC) and EvalPlus (HumanEval).
To demonstrate Code2Bench’s effectiveness in evaluating LLMs on real-world code complexities and revealing specific failure modes often missed by benchmarks based on simpler tasks, we compare the performance of state-of-the-art models on Code2Bench-2505 (Python SC subset) against EvalPlus’s evaluation on HumanEval [21, 14]. Both benchmarks strive for rigorous functional verification, with EvalPlus augmenting HumanEval’s test suite using mutation-based testing and Code2Bench employing Property-Based Testing (PBT) on tasks derived directly from recent real-world code. By comparing performance on these two distinct benchmarks, we can gain crucial insights into how task source (real-world dynamic vs. manually-crafted static) and test generation methodology (PBT for complex/constrained structures vs. mutation for variations) impact the revealed model capabilities and limitations. Figure 3 presents a scatter plot visualizing the Pass@1 results for a range of models evaluated on both Code2Bench-2505 (Python SC) and EvalPlus (HumanEval).
Figure 3 reveals a significant performance gap between Code2Bench-2505 and EvalPlus (HumanEval), with models consistently achieving substantially lower Pass@1 scores on Code2Bench-2505. This indicates that tasks derived from real-world code complexities in Code2Bench-2505 are generally more challenging for current LLMs compared to HumanEval’s tasks, even when both benchmarks employ rigorous testing. The scatter plot also shows distinct evaluation profiles; while a general correlation exists, models exhibit varied relative performance, suggesting the benchmarks evaluate different facets of code generation ability. This performance disparity and the differing evaluation profiles demonstrate that Code2Bench, by sourcing tasks from the dynamic and nuanced landscape of real-world software development and applying rigorous PBT tailored for complex inputs, effectively exposes LLM limitations and failure modes specific to real-world code complexities that may not be as apparent on benchmarks based on simpler, static tasks like HumanEval.
4.6 Performance by Difficulty Levels
This subsection analyzes model performance on Code2Bench-2505 tasks based on their classified difficulty levels (Easy, Medium, Hard), providing insight into how LLM capabilities scale with task complexity. We examine the average Pass@1 performance across all evaluated models SC and WSC Python tasks across these difficulty levels. The results are summarized in Table 4. More detailed results are available in Appendix I.3. Table 4 reveals a clear trend: for both SC and WSC Python
Table 4: Average Pass@1 by Difficulty in Python.
Task Type Easy(%) Medium(%) Hard(%)
SC 41.94 28.42 28.03
WSC 65.75 36.11 15.12
tasks, average Pass@1 performance decreases significantly with increasing difficulty. Easy tasks show the highest performance, followed by Medium, with Hard tasks exhibiting substantially lower Pass@1 in both categories. This confirms that Code2Bench-2505 successfully captures a wide spectrum of real-world coding challenges, from relatively straightforward problems to highly difficult ones. While WSC tasks generally show higher Pass@1 than SC tasks across Easy and Medium levels, the performance gap narrows or reverses at the Hard level, where both types of complex tasks pose significant challenges to current LLMs. This pronounced drop in performance at higher difficulty levels highlights Code2Bench’s ability to generate challenging tasks that effectively differentiate model capabilities and expose their limitations on complex real-world coding problems.
@@ 4.7 The Effectiveness of PBT-Generated Tests
To further demonstrate the rigor and discriminative power of the PBT-generated test suites in Code2Bench-2505, we analyzed the distribution of test case pass rates. While Pass@1 (100% passing) is our primary metric, examining the percentage of problems where models achieved a test case pass rate $\geq 98\%$ (indicating $\leq 10$ failures out of $\approx 500$ tests) provides crucial insight into the capture of subtle errors. Table 5 presents this percentage for four representative models across various task types (SC tasks in Python, Go, Java, TS, JS, and WSC tasks in Python). The metric is computed as the count of problems passing $\geq 98\%$ of tests ($\leq 2\%$ failed) divided by the total number of distinct problems in each respective category.
Table 5: Percentage of Problems with Test Case Pass Rate $\geq$ 98% for Representative Models (%).
Model Name SC: Python SC: Go SC: Java SC: TS SC: JS WSC: Python
Claude-3.7-sonnet 13.5 14.7 15.6 10.9 18.6 8.8
Gemini-2.5-flash-preview 11.9 11.8 9.5 10.4 19.1 9.8
DeepSeek-V3 11.9 16.1 16.5 10.9 20.9 7.8
Qwen3-235b-a22b 12.9 10.8 10.4 10.4 19.5 7.3
As shown in Table 5, even for these capable models, the percentage of problems with a test case pass rate of 98% or higher generally ranging from approximately 7% to 21%. This highlights that models frequently produce code that is nearly correct but fails on a small number of test cases. If these specific, subtle-error-revealing test cases were absent, these problems would be incorrectly marked as perfectly solved. This underscores the effectiveness of our comprehensive PBT-generated test suites in capturing nuanced failures on complex, real-world code, providing a more reliable assessment than possible with less thorough testing. More details can be found in Appendix I.4.
@@ 5 Related Work
Evaluating Large Language Models (LLMs) on code generation tasks is an active research area, with numerous benchmarks proposed. Early benchmarks[42, 43] like HumanEval [14], MBPP [15], and APPS [16] provide static collections of isolated code snippets with tests, proving valuable for initial model development but facing limitations regarding dataset contamination [10, 11, 12, 13] and the lack of real-world context and dependencies. Efforts like EvalPlus [21] enhance static benchmarks with more robust tests via mutation. However, these may not fully represent real-world code complexities or provide automated construction from code repositories. Multilingual benchmarks[44] like HumanEval-X [28] and Aider polyglot benchmark [45] evaluate cross-language abilities but are at risk of data leakage. Benchmarks focusing on repository-level context or dependencies[46, 20, 47, 48] include RepoBench [24], CrossCodeEval [25], R2E [35], DevEval [26], BigCodeBench [23], and CODEAGENT [27]. WebBench [49] introduces sequential, real-world web development tasks. While these capture aspects of real-world interaction, they often lack sufficient testing. Code2Bench stands out by offering an end-to-end pipeline for dynamically generating rigorous benchmark instances from recent real-world GitHub repositories.
@@
A key limitation of static benchmarks is the risk of test set contamination, where models may have been trained on the same data used for evaluation[10, 11, 12, 13]. This has motivated the development of "live" benchmarks. DynaBench [50] identified these challenges and advocated for continuously evolving benchmarks. Chatbot Arena [17] provides a platform for dynamic evaluation based on user interactions. LiveBench [18] sources new data from specific, non-code domains like mathematics and news, while LiveCodeBench [19] collects recent problems from competitive programming platforms. Other methods leverage LLMs for task mutation to generate new problems, such as EvoEval [51]. Evocodebench [20] introduces a periodically updated benchmark to mitigate leakage, but currently focuses on Python and may lack rigorous test. Arena-Hard-Auto [52] filters crowdsourced prompts into benchmarks. While existing initiatives have made progress in addressing contamination and enabling dynamic evaluation, they often rely on specific data sources, focus on evaluation platforms, or lack a systematic, automated pipeline for constructing high-quality benchmarks from real-world code at scale. Code2Bench bridges this gap by providing a novel, automated, end-to-end pipeline that dynamically extracts, filters, and constructs rigorous benchmark instances from recent activity in real-world GitHub repositories.

### v2

@@ 4.1 Experimental Setup
Evaluated Models. We selected a diverse suite of 10 state-of-the-art Large Language Models, encompassing both leading closed-source APIs and prominent open-source families. A cornerstone of our evaluation integrity is the strict prevention of data contamination. The Code2Bench-2509 benchmark was constructed exclusively from code committed after May 2025, a date subsequent to the knowledge cutoff of all evaluated models.
Evaluation Protocol. Our primary metric is Pass@1 Kulal et al. (2019), which measures the functional correctness of the first-generated solution and closely mirrors a developer’s real-world experience with coding assistants Jain et al. (2024a). All evaluations were conducted in a zero-shot, deterministic setting, employing greedy decoding (temperature 0) as is standard practice Zhuo et al. (2024); Roziere et al. (2023). For each task, the model received a standardized instruction containing the function signature and a natural language description(See more details in Appendix I.2).
Execution Environment. The generated code for each task was executed in a sandboxed environment against the full suite of PBT-generated tests (5̃00 per task). A language-specific test runner performed differential testing, comparing the output of the model-generated code against the ground-truth implementation. A task is considered passed only if it correctly solves all test cases. Open-source models were served via vLLM Kwon et al. (2023), while others were accessed through their official APIs.
4.2 A Multi-Dimensional Diagnostic of LLM Capabilities
Table 3: Pass@1 performance (%) on the CODE2BENCH-2509 suite.
Model SC-Python WSC-Python SC-Java
(%) [95% CI] (%) [95% CI] (%) [95% CI]
Closed-Source Models
Claude-4-sonnet 40.1 [33.6 – 46.5] 38.7 [32.0 – 45.4] 47.4 [40.9 – 53.4]
Gemini-2.5-Flash 37.8 [30.9 – 44.2] 36.6 [29.4 – 43.3] 45.0 [39.0 – 51.0]
Open-Source Models (Ordered by Scale)
DeepSeek-V3 34.4 [28.4 – 40.4] 37.6 [31.4 – 44.3] 47.8 [41.4 – 54.2]
Qwen3-235b-a22b 34.6 [28.6 – 41.0] 36.6 [29.9 – 43.3] 46.6 [40.9 – 53.0]
Llama-4-scout 25.8 [19.8 – 31.8] 32.5 [26.3 – 39.2] 44.2 [37.8 – 49.8]
Qwen3-32b 31.3 [25.8 – 36.9] 34.5 [27.8 – 41.2] 43.0 [37.4 – 49.4]
Mistral-small-3.1 (24B) 30.4 [24.4 – 36.9] 38.7 [32.5 – 45.9] 43.4 [37.4 – 49.4]
Qwen3-8b 25.1 [19.3 – 31.4] 34.0 [27.8 – 40.7] 39.0 [32.9 – 44.2]
Gemma-3n-e4b-it 22.6 [17.1 – 28.6] 26.3 [19.6 – 32.5] 34.5 [28.5 – 40.2]
Qwen3-1.7b 14.3 [9.7 – 19.4] 16.5 [11.3 – 21.6] 17.7 [12.8 – 22.5]
Figure 3: Fingerprints across the three evaluation tracks—SC-Python (left), WSC-Python (middle), and SC-Java (right)—shown as ridgeline plots. Each curve captures a model’s outcome distribution, ranging from SyntaxErr to Perfect, with key pass rates annotated.
A primary limitation of existing benchmarks is their inability to provide deep, diagnostic insights. We argue this stems from two fundamental constraints: a narrow source of problems and superficial testing. The CODE2BENCH framework overcomes these limitations through two core principles: scaling the source from dynamic, real-world code, and scaling the rigor of evaluation via Property-Based Testing (PBT). In this section, we demonstrate how this dual-scaling approach enables an unprecedentedly fine-grained diagnostic of LLM coding capabilities.
Table 3 presents the overall Pass@1 performance, revealing clear capability tiers among models. The data indicates that performance varies significantly across our three benchmark components, hinting at the different skills they evaluate. To move beyond these aggregate scores and understand the underlying reasons for these variations, we now turn to a more fine-grained analysis.
To dissect how and why models succeed or fail, we introduce a granular, multi-stage outcome spectrum, visualized as a diagnostic fingerprint for each model in Figure 3. Outcome categories are ordered to reflect a progression from catastrophic failure to complete success: SyntaxErr (code fails to compile/run), RuntimeErr (code crashes during execution), LogicErr (code runs but is logically incorrect on all test cases), followed by a four-tier partial success range based on pass rates—from Fail (<20%) to NearPerf (>=98%)—and culminating in Perfect solutions. Figure 3 visualizes each model’s distribution across this spectrum. The synergy between the aggregate scores in the table and these distributional fingerprints reveals two profound insights into the nature of LLM coding intelligence.
Decoupling Algorithmic Synthesis from API Application.
The diagnostic fingerprints (Figure 3, left vs. middle) show a systematic shift in failure modes: on SC-Python, the dominant failure is LogicErr, indicating a core challenge in first-principle reasoning. Conversely, on WSC-Python, this peak vanishes and RuntimeErr emerges as a primary obstacle, suggesting the challenge shifts to the correct application of external APIs.
Language Paradigms as Performance Scaffolding.
The comparison between SC-Python and SC-Java (Figure 3, left vs. right), reveals the profound impact of language paradigms. In Java, the prominent LogicErr and RuntimeErr peaks seen in Python are sharply suppressed, while performance in the Perfect category surges for all models. We hypothesize this is not because models are inherently "better" at Java, but because its static type system acts as a powerful “performance scaffolding”, pruning a vast space of potential errors at compile time. This demonstrates that an LLM’s coding ability is not an abstract quantity but is fundamentally intertwined with the target language’s ecosystem, a crucial interaction our framework is the first to systematically quantify.
@@ 4.3 The Effectiveness of PBT-Generated Tests
A core tenet of the Code2Bench framework is scaling test rigor via Property-Based Testing (PBT). This section provides quantitative evidence for the necessity of this approach by analyzing the prevalence of "Near-Perfect" failures—solutions that pass at least 98% of our comprehensive test suite but fail on a handful of subtle edge cases. These instances represent an “illusion of correctness” that would likely go undetected by conventional, less rigorous benchmarks.
Table 4 quantifies the frequency of these "near-miss" failures, where a solution passes the vast majority of test cases (>98%) but ultimately fails. The data reveals a significant and consistent pattern: on average, 6.94% of submissions for SC-Python tasks fall into this treacherous category.
This finding directly underscores the critical necessity of our Property-Based Testing (PBT) methodology. Without the exhaustive, edge-case-driven verification enabled by PBT, these nearly 7% of submissions—which are functionally almost correct—would have been falsely classified as successes by conventional, sparse test suites. This would lead to a significant overestimation of model capabilities.
Table 4: Prevalence of “Near-Perfect” Failures (Pass@ $\geq$98%) in CODE2BENCH.
Model SC-Py WSC-Py SC-Java
Claude-Sonnet-4 8.76 5.15 2.41
Gemini-2.5-Flash 6.45 5.67 4.02
DeepSeek-V3 7.80 3.61 2.41
Qwen3-235b-a22b 6.45 3.61 2.01
LLAMA-4-scout 8.29 5.15 2.01
Mistral-Small-3.1 6.91 5.67 2.01
Qwen3-32b 7.37 3.61 1.61
Qwen3-8b 6.76 4.64 2.41
Gemma-3n-e4b-it 5.53 5.67 2.81
Qwen3-1.7b 5.07 3.09 0.80
Avg. (%) 6.94 4.57 2.25
Top-performing models are not immune to this illusion of correctness; “DeepSeek-V3” and “Claude-4-sonnet”, for example, see approximately 8% of their submissions fall into this category. This demonstrates that even the most capable models consistently struggle with the final frontier of logical robustness, a weakness that only a truly rigorous testing paradigm like PBT can reliably expose.
Interestingly, the rate of near-perfect failures is lower in WSC-Python (4.57%) and lowest in SC-Java (2.25%). This aligns with our findings in Section 4.2. In WSC tasks, the problem is often a binary choice of the correct API call, leaving less room for "almost correct" logic. In Java, the strict type system likely prevents many of these subtle logical errors at the compilation stage. Therefore, the high rate of near-perfect failures in SC-Python highlights its unique position as the most challenging testbed for a model’s pure, unaided logical robustness.
Ultimately, this analysis validates the central role of rigorous, scaled testing. By systematically uncovering these near-perfect failures, Code2Bench provides a more accurate measure of a model’s true capabilities and offers invaluable, fine-grained feedback for identifying and rectifying their most subtle weaknesses.
4.4 The Impact of Dynamic Sourcing and Real-World Complexity
To situate Code2Bench-2509 within the existing landscape, we conduct a direct comparison against EvalPlus, a state-of-the-art benchmark that enhances HumanEval/MBPP with more rigorous, mutation-based testing. While EvalPlus represents the pinnacle of evaluation on static, well-known problem sets, Code2Bench-2509 introduces the dimensions of dynamic sourcing and real-world complexity. This comparison aims to answer a critical question: how does a model’s performance on canonical programming puzzles translate to its ability to handle fresh, complex code from the wild?
Figure 4: Performance on Evalplus and Code2Bench-2509
The head-to-head comparison is visualized in Figure 4, which plots the Pass@1 scores of ten prominent LLMs on our SC-Python-2509 (X-axis) against their performance on HumanEval (Y-axis). The results are stark and reveal three critical insights:
A Systematically More Challenging Benchmark. The most striking observation is that all models are located deep in the red-shaded region, far above the $y=x$ diagonal of equal performance. This demonstrates that Code2Bench-2509 presents a systematically higher level of difficulty for all models, without exception. For instance, top-performing models like Claude-4-Sonnet, which achieve a near-perfect score of 97% on HumanEval, see their performance plummet to 40.1% on our benchmark—a drop of over 50 percentage points. This substantial performance gap suggests that high scores on legacy benchmarks may create an illusion of capability that does not hold up against the complexity of novel, real-world code.
Probing Generalization Over Memorization. This performance delta is not merely a matter of difficulty, but of the fundamental capabilities being measured. HumanEval’s problems are years old and likely part of the training corpora. The lower performance on SC-Python-2509—a benchmark guaranteed to be unseen—strongly indicates that our evaluation measures a model’s true generalization ability on novel algorithmic challenges, rather than its capacity for pattern memorization. This underscores the critical need for dynamic, contamination-resistant benchmarks to ensure a fair and realistic assessment of an LLM’s problem-solving intelligence.
@@ 5 Related Work
Evaluating Large Language Models (LLMs) on code generation tasks is an active research area, with numerous benchmarks proposed. Early benchmarksCassano et al. (2023); Li et al. (2022) like HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and APPS Hendrycks et al. (2021) provide static collections of isolated code snippets with tests, proving valuable for initial model development but facing limitations regarding dataset contamination Carlini et al. (2021); Sainz et al. (2023); Yang et al. (2023); Team et al. (2024) and the lack of real-world context and dependencies. Efforts like EvalPlus Liu et al. (2023a) enhance static benchmarks with more robust tests via mutation. However, these may not fully represent real-world code complexities or provide automated construction from code repositories. Multilingual benchmarksYan et al. (2023) like HumanEval-X Zheng et al. (2023b), Aider polyglot benchmark 4 and AutoCodeBench Chou et al. (2025) evaluate cross-language abilities but are at risk of data leakage. Benchmarks focusing on repository-level context or dependenciesTang et al. (2023); Li et al. (2024b); Yu et al. (2024); Wang et al. (2024) include RepoBench Liu et al. (2023b), CrossCodeEval Ding et al. (2023), R2E Jain et al. (2024b), DevEval Li et al. (2024c), BigCodeBench Zhuo et al. (2024), and CODEAGENT Zhang et al. (2024). WebBench Xu et al. (2025) introduces sequential, real-world web development tasks. While these capture aspects of real-world interaction, they often lack sufficient testing. Code2Bench stands out by offering an end-to-end pipeline for dynamically generating rigorous benchmark instances from recent real-world GitHub repositories.
@@
A key limitation of static benchmarks is the risk of test set contamination, where models may have been trained on the same data used for evaluationCarlini et al. (2021); Sainz et al. (2023); Yang et al. (2023); Team et al. (2024). This has motivated the development of "live" benchmarks. DynaBench Kiela et al. (2021) identified these challenges and advocated for continuously evolving benchmarks. Chatbot Arena Chiang et al. (2024) provides a platform for dynamic evaluation based on user interactions. LiveBench White et al. (2024) sources new data from specific, non-code domains like mathematics and news, while LiveCodeBench Jain et al. (2024a) collects recent problems from competitive programming platforms. Other methods leverage LLMs for task mutation to generate new problems, such as EvoEval Xia et al. (2024). Evocodebench Li et al. (2024b) introduces a periodically updated benchmark to mitigate leakage, while DOMAINEVAL Zhu et al. (2025) employs dynamic data sources with automated updates for the same purpose. Both efforts currently focus on Python and may lack rigorous test. Arena-Hard-Auto Li et al. (2024d) filters crowdsourced prompts into benchmarks. While existing initiatives have made progress in addressing contamination and enabling dynamic evaluation, they often rely on specific data sources, focus on evaluation platforms, or lack a systematic, automated pipeline for constructing high-quality benchmarks from real-world code at scale. Code2Bench bridges this gap by providing a novel, automated, end-to-end pipeline that dynamically extracts, filters, and constructs rigorous benchmark.